Grant had 8 1/5 kilograms of flour and used 3/7 of it. How much flour did grant use?

Found 2 solutions by stanbon, mallikarjuna1st:
Answer by stanbon(75887)

8 1/5 = 41/5
(41/5)(3/7)]=123/35 kilograms
Cheers,
Stan H.

Answer by mallikarjuna1st(3)

41/5 * 3/7 = 123/35